What is the domain of the composite function G(F(x))?
fx=√x
G(F(x))=1/√x-4
G(y)=1/y-4
the possible awnsers are
A. x < 0
B. x ≤ 4
C. x ≥ 0
D. x > 0
E. x < 4
F. All real numbers


Answer by stanbon(75887)

What is the domain of the composite function G(F(x))?
f(x) = √x
G(F(x))=1/√x-4
G(y)=1/y-4
the possible answers are
A. x < 0
B. x ≤ 4
C. x ≥ 0
D. x > 0
E. x < 4
F. All real numbers
-------------------------
f(x) = √x
Domain of f is x>=0
--------------------
G(y)=1/y-4
Domain of G eliminates 4 from the domain
---------------
G(F(x))=1/√x-4
Domain of the composite is "All Real numbers >=0 except x=4
==================
Cheers,
Stan H.